# Hierarchical Hybrid Learning for Long-Horizon Contact-Rich Robotic Assembly

### v1

@@ Abstract
Generalizable long-horizon robotic assembly requires reasoning at multiple levels of abstraction. End-to-end imitation learning (IL) has been proven a promising approach, but it requires a large amount of demonstration data for training and often fails to meet the high-precision requirement of assembly tasks. Reinforcement Learning (RL) approaches have succeeded in high-precision assembly tasks, but suffer from sample inefficiency and hence, are less competent at long-horizon tasks. To address these challenges, we propose a hierarchical modular approach, named ARCH (Adaptive Robotic Compositional Hierarchy), which enables long-horizon high-precision assembly in contact-rich settings. ARCH employs a hierarchical planning framework, including a low-level primitive library of continuously parameterized skills and a high-level policy. The low-level primitive library includes essential skills for assembly tasks, such as grasping and inserting. These primitives consist of both RL and model-based controllers. The high-level policy, learned via imitation learning from a handful of demonstrations, selects the appropriate primitive skills and instantiates them with continuous input parameters. We extensively evaluate our approach on a real robot manipulation platform. We show that while trained on a single task, ARCH generalizes well to unseen tasks and outperforms baseline methods in terms of success rate and data efficiency. Videos and code can be found at https://long-horizon-assembly.github.io.
I Introduction
@@
Figure 1: a) A Zivid camera captures RGB-D images while a UR10e robot equipped with force-torque sensors is used for the manipulation tasks. b) Objects are randomly placed within a region on the table. The robot must insert them into the correct receptacle in the Insertion Board, adjusting their orientation with the Fixture if necessary. c) The library of RL and model-based action primitives.
The manufacturing industry is increasingly turning to robotic systems for complex assembly tasks, driven by the need for greater precision, consistency, and efficiency [12, 10]. Nevertheless, long-horizon, contact-rich, and high-precision robotic assembly tasks continue to pose significant challenges in robotics and automation, as these tasks demand sophisticated learning and object interaction capabilities that go beyond traditional programming approaches [17]. Our research endeavors to ground the robotic assembly to real-world applications where the robots are asked to perform intricate and long-horizon assembly tasks widely demanded in industrial environments, e.g., parts insertion [28] and cable routing [27]. The acceleration of complex assembly tasks is a crucial step toward more flexible and adaptive industrial processes. By focusing on automating such long-horizon tasks, we seek to enhance productivity and efficiency across various sectors, including manufacturing, construction, and logistics.
@@
Current industrial robotics applications to assembly and manufacturing are largely engineered for a specific task and struggle with adapting to varied assembly scenarios and component variations. Ideally, such systems are capable of handling high-precision contact-rich operations under uncertainty  [20] and generalizing to new scenarios from limited demonstrations  [47] using multimodal inputs such as visual and force-torque feedback [35, 53].
@@
Recent advances in learning-based solutions have started to tackle some of these problems. End-to-end imitation learning (IL) from human demonstrations has made marked progress [8], but still fails to generalize outside a narrow training distribution, requires significant data from a trained expert demonstrator  [18], and often struggles with high-precision tasks. Simply sequencing individual primitives can lead to compounding errors, diminishing overall system performance [42]. While reinforcement learning (RL) can drive specific assembly operations, it usually struggles with more complicated long-horizon tasks [41], where the training faces challenges of sparse rewards and an exploding sampling space due to larger long-horizon footprints.
To address these challenges, we propose a hierarchical approach for robotic assembly, especially targeted at long-horizon, contact-rich, high-precision settings. Our method employs a hierarchical strategy that involves developing a set of low-level primitive skills (e.g. grasp, insert) and a high-level policy to select and integrate these primitives. The framework combines classic motion planning algorithms with RL policies to develop low-level primitive skills, achieving both efficiency and adaptability. More specifically, we leverage motion planning algorithms [5] to efficiently guide the end-effector to target goals. For complex assembly operations in partially observed environments, we utilize RL policies that enable adaptive, contact-rich, and high-precision manipulation. The RL policies are trained fully in simulation with domain randomization such that they can directly transfer to the physical world without any fine-tuning. Based on these primitive skills, the high-level policy is learned from a few trajectories of human demonstrations to select and organize among these skills to achieve the final assembly goal. This hybrid strategy allows our system to leverage pre-existing knowledge for routine movements while maintaining the flexibility to learn and adapt to new, intricate assembly tasks in unfamiliar settings. Moreover, our high-level primitive selection policy has a small action space and is object-agnostic. Hence, it can be effectively trained via imitation learning, as it only requires a few human demonstrations.
@@
  • We introduce ARCH, a hierarchical framework designed to tackle the challenging problem of long-horizon robotic assembly.
  • Our hierarchical framework includes a low-level skill library and a high-level imitation-learned policy that selects and composes the primitive skills. The low-level skills are built upon efficient motion planning algorithms for end-effector movement and RL policies for high-precision, adaptive assembly in contact-rich scenarios.
  • We conducted extensive hardware experiments on a robot manipulator in an assembly work cell. Despite being trained on a single task, our approach archives generalization to other novel tasks and outperforms baseline methods for long-horizon robotic assembly tasks.
II Related Work
Figure 2: We propose a hierarchical framework for long-horizon robotic assembly. The high-level policy $\pi_{\theta}$, which takes as input object pose from pose estimation and robot proprioception, and outputs a categorical distribution to select the best low-level primitive as well as its continuous parameters, is obtained via imitation learning. The low-level policy executes the selected primitive using either an RL-based or a motion-planned (MP) policy. We train in simulation an RL policy for the contact-rich portion of the task, e.g. insertion, based on force-torque feedback. We use the MP policies for primitives that are in free space, e.g. move.
II-A Task and Motion Planning
@@
Task and Motion Planning (TAMP) [21, 22, 16] is an effective approach for long-horizon manipulation problems as it can resolve temporally dependent constraints through hybrid symbolic-continuous reasoning [40, 19, 4, 44]. These plans may involve regrasping [1], clearing obstructing objects [13], or moving to gather information [14]. Despite these abilities, such methods typically require hand-designed symbolic transition functions and continuous parameter samplers. Additionally, TAMP methods are computationally expensive, which limits their ability to handle failures in dynamic tasks. We argue that the symbolic transition functions and samplers can be replaced by a learnable high-level policy to address the issues mentioned above.
II-B Learning for Long-Horizon Manipulation
@@
There is a large body of research in learning for long-horizon manipulation tasks. Diffusion policies [8, 9, 52, 45] have shown to be a powerful tool for addressing complex manipulation tasks by leveraging their generative and multi-modality capabilities. However, they require a large amount of training data, i.e. human demonstrations, to learn effective end-to-end policies for complex, long-horizon tasks [6]. Therefore, a significant amount of research has adopted the divide-and-conquer mindset and tackled the task by decomposing it into easier and reusable subtasks, instead of learning an entire task with a single policy.
@@
Skill decomposition and chaining [24, 23, 2, 25, 26, 11] is a promising way to synthesize long-horizon and complex behaviors by sequentially chaining previously learned simpler skills through transition functions. Mishra et al. [33] trains individual skill diffusion models as action primitives and combine them at test time, when learned distributions of the skills are linearly chained to solve for a long-horizon goal during evaluation. Mao et al. [31] learns primitive skills from human demonstrations based on haptic feedback and segments a long-horizon task into a sequence of skills for dexterous manipulation. Chen et al. [7] introduces a bi-directional optimization framework that chains RL-trained sub-policies together using a transition feasibility function for long-horizon dexterous manipulation.
II-C Hierarchical Modeling in Robotics
@@
Hierarchical approaches typically offer policies at varying abstraction levels [39]. Xian et al. [48] and Ma et al. [29] propose to learn a high-level policy from demonstrations to predict end-effector key poses and a low-level diffusion-based trajectory generator for connecting these key poses to achieve the final goal. MimicPlay [47] learns a high-level plan from human videos of manipulating objects and low-level visuomotor controls from teleoperated demonstrations on the real robot. MAPLE [36] enhances standard RL algorithms by incorporating a predefined library of behavioral primitives. The most relevant work to our approach uses hierarchical policy decomposition for a multi-stage cable routing task [27] . They define scripted and imitation learned low-level primitives and learn a high-level policy to select which primitive to execute from human demonstration, all in real. We combine model-based and model-free primitives on the low level and learn a high-level policy for parameterized primitive selection and instantiation. We also train a robust RL policy for the contact-rich primitive - insertion - in simulation and perform zero-shot sim-to-real transfer. Additionally, we train all primitive modules in advance, and allow expert demonstrators to make use of these primitives as necessary.
Recent advancements in robotic learning have also highlighted the need for benchmarks that effectively balance generalization and the complexity of manipulation tasks. The Functional Manipulation Benchmark (FMB) [28] proposed by Luo et al. addresses this gap by defining functional manipulation as a series of relevant behaviors, such as grasping, repositioning, and physical interaction with objects. This benchmark emphasizes the importance of both contact dynamics and object generalization, making it a valuable resource for researchers aiming to develop robots capable of performing intricate assembly tasks. For these reasons, we focus on the multi-peg assembly benchmark, introduced in FMB [28], to showcase the generalizability of our approach.
III Background
@@
In this work, we model the long-horizon robot assembly task as a parameterized-action Markov decision process (PAMDP) [32]. A PAMDP problem consists of the tuple $\langle\mathcal{S},\mathcal{A},P,R,\gamma\rangle$ where $\mathcal{S}$ is the continuous state space, $\mathcal{A}$ is a set of discrete action primitives $\{a_{1},a_{2},...,a_{k}\}$, each with $m_{a}$ continuous parameters $X_{a}\subseteq{\mathbb{R}^{m_{a}}}$, $P(s,a,s^{\prime})$ is the probability of transitioning to state $s^{\prime}$ when taking action primitive $a$ in state $s$, $R(s,a)$ is the reward from taking action primitive $a$ in state $s$, and $\gamma$ is the discount factor. Our goal in a PAMDP is to find a policy $\pi(a,x|s)$ that minimizes the discrepancy between the learned policy and the expert demonstrations that maximize reward under the PAMDP.
IV Method
Figure 3: High-Level Policy Architecture.
@@ IV Method
Solving a PAMDP involves not only selecting the best discrete primitive but also optimizing over the continuous parameter space for each primitive at each state. In this work, we fix a set of parameterized primitives that are building blocks of manipulation for assembly tasks (See Section IV-A) and attempt to learn the high-level policy $\pi_{\theta}(a,x|s)$ as a neural network with parameters $\theta$. This results in a hierarchical framework, shown in Fig. 2. In this section, we describe our low-level primitive library, high-level policy architecture, data collection pipeline, and the pose estimation module.
@@
IV-A Low-level Primitive Library: Modeling Basic Skills with Motion Planning and Reinforcement Learning
@@
Our focus is on providing agents with a library of flexible primitives that act as foundational components for high-precision, contact-rich robotic assembly tasks. The hierarchical decision-making system operates independently of the specific implementations of these primitives, which may include either closed-loop, learning-based skills or analytical motion planners. Regardless of how they function internally, it is essential that the primitives are adaptable to varying behaviors, which is why we introduce parameters $x$ to customize each primitive. In our learning framework, we treat these primitives as functional APIs, where input parameters $x$ define the actions to be executed. These input parameters typically have clear semantics, such as specifying a 6-DoF end-effector pose for a grasping primitive. While these primitives offer flexibility, we acknowledge that they may not cover all possible scenarios, and relying solely on them could restrict the range of behaviors an agent can achieve.
We design our low-level primitive library according to the building blocks of manipulation for assembly tasks. These primitives are not exhaustive and easily extendable.
Grasp: The robot moves its end-effector to a pre-grasp pose $p_{g}\in\mathbf{SE}(3)$, specified by the input parameter, and closes its gripper. A motion planner is used for its execution.
Place: The robot moves its end-effector to a pre-place pose $p_{p}\in\mathbf{SE}(3)$, specified by the input parameter, and opens its gripper. A motion planner is used for its execution.
Move: The robot moves its end-effector to a pose $p\in\mathbf{SE}(3)$, specified by the input parameter. A motion planner is used for its execution.
@@
Insert: The robot has the object in grasp and inserts it into the correct hole on the insertion board. An RL policy $\pi_{L}(\phi)$ is used because the insertion task involves rich contact between the object and its receptacle. We train an RL policy in simulation [34] using PPO [38]. Its observation space includes end-effector (EE) force-torque (FT) and EE pose relative to the target pose; its action is EE velocity $u_{t}$; and its reward is the negative distance between the current pose $s^{p}_{t}$ and the target pose $g$.
$\tilde{r}(s^{p}_{t},u_{t},g)=-||s^{p}_{t}-g||_{2},$ (1)
@@
The objective of goal-conditioned Insertion primitive is to reach goal pose $g$ that maximize the expectation of the cumulative return as Eq. 2 shows:
@@
IV-B High-level Policy: Composing Primitives via Imitation
The high-level policy $\pi_{\theta}(a,x|s)$ takes object pose obtained from pose estimation (see Section IV-B2) and robot proprioception as inputs $s$, to select the appropriate primitive $a$ from the low-level primitive library and predict its continuous parameter $x\in\mathbb{R}^{d_{\textrm{control}}}$ for low-level control. We collect human demonstration data (see Section IV-B1) and train the high-level policy via imitation learning. As shown in Fig. 3, we modify the Diffusion Transformer (DiT) [37] architecture as the backbone of our high-level policy, considering its strong sequential data handling capability is suitable for handling with a history of previous states and actions. The DiT outputs softmax scores for each primitive and the continuous action parameter for the primitive with the highest score. These two functions are supervised using cross-entropy loss and MSE loss, respectively. We employ DiT blocks with adaptive LayerNorm-Zero conditioning. We use the Robomimic observation encoder [30] to extract features from object pose ($s^{op}$) and proprioceptive pose ($s^{p}$). The goal of imitation learning is to find a parameterized policy $\pi_{\theta}$ that can maximize the likelihood function based on the currently collected demonstration data $\mathcal{D}=\{(\mathbf{s,a,x)}\}$:

### v2

# ARCH: Hierarchical Hybrid Learning for Long-Horizon Contact-Rich Robotic Assembly

@@ Abstract
Generalizable long-horizon robotic assembly requires reasoning at multiple levels of abstraction. While end-to-end imitation learning (IL) is a promising approach, it typically requires large amounts of expert demonstration data and often struggles to achieve the high precision demanded by assembly tasks. Reinforcement learning (RL) approaches, on the other hand, have shown some success in high-precision assembly, but suffer from sample inefficiency, which limits their effectiveness in long-horizon tasks. To address these challenges, we propose a hierarchical modular approach, named Adaptive Robotic Compositional Hierarchy (ARCH), which enables long-horizon, high-precision robotic assembly in contact-rich settings. ARCH employs a hierarchical planning framework, including a low-level primitive library of parameterized skills and a high-level policy. The low-level primitive library includes essential skills for assembly tasks, such as grasping and inserting. These primitives consist of both RL and model-based policies. The high-level policy, learned via IL from a handful of demonstrations, without the need for teleoperation, selects the appropriate primitive skills. We extensively evaluate our approach in simulation and on a real robotic manipulation platform. We show that ARCH generalizes well to unseen objects and outperforms baseline methods in terms of success rate and data efficiency. More details are available at: https://long-horizon-assembly.github.io.
Keywords: Robotic Assembly, Long-horizon Assembly, Hybrid Learning
1 Introduction
@@
Figure 1: a) Workcell setup: a camera captures RGB-D images while a robot equipped with a force-torque sensor is used for the manipulation tasks. b) Task example: objects are randomly placed within the Grasp Region on the table. The robot must insert them into the correct receptacle in the Insertion Board, adjusting their orientation with the Fixture if necessary. c) Examples of the library of RL and model-based action primitives. d) Evaluation tasks: FMB Assembly, 5-part Beam Assembly, and 9-part Stool Assembly.
@@
The manufacturing industry is increasingly turning to robotic systems for assembly tasks, driven by the need for greater precision, consistency, and efficiency [1, 2]. Nevertheless, long-horizon, contact-rich, and high-precision robotic assembly tasks continue to pose significant challenges in robotic automation, as these tasks demand sophisticated learning and object interaction capabilities that go beyond traditional approaches [3]. Current industrial robotic applications to assembly and manufacturing are largely engineered for a specific task and struggle with adapting to varied assembly scenarios and component variations. Ideally, such systems should be capable of handling high-precision contact-rich operations [4] and generalizing to long-horizon new scenarios from limited demonstrations  [5, 6] using multimodal inputs such as visual and force-torque feedback [7, 8].
@@
Toward this goal, recent advances in learning-based solutions have started to tackle some of these problems. End-to-end imitation learning (IL) from human demonstrations has made encouraging progress [9, 10, 11], but usually requires a significant amount of expert demonstration data [12], which can be costly to collect, and often struggles with high-precision tasks. While reinforcement learning (RL) can drive specific assembly operations, such as high-precision insertion, it usually falls short on more complicated long-horizon tasks [13, 14, 15, 16, 17], where training faces challenges of sparse rewards and an exploding sampling space due to the larger long-horizon footprint. Alternatively, naively sequencing atomic actions or behaviors can lead to compounding errors, diminishing overall system performance [18].
To address these challenges, we propose a hierarchical approach for robotic assembly, especially targeted at long-horizon, contact-rich, high-precision settings. Our method adopts a hierarchical framework that consists of a library of parameterized low-level primitive skills (e.g., grasp, insert) and a high-level policy that sequences these primitives based on task context. To build robust and adaptable low-level primitives, we leverage both classical motion planning (MP) algorithms and RL policies. This hybrid MP-RL strategy enables the system to leverage prior knowledge for efficient execution of routine motions while remaining flexible enough to handle contact-rich and uncertain dynamics. The RL components are trained in simulation using domain randomization, enabling feasible transfer to real-world execution. At the high level, the policy operates over a compact action space, enabling efficient learning from a handful of human demonstrations via imitation learning, thereby alleviating the data collection burden common in end-to-end imitation learning pipelines.
@@
  • We introduce Adaptive Robotic Compositional Hierarchy (ARCH), a hierarchical framework designed to tackle the challenging problem of long-horizon robotic assembly.
  • We develop a hybrid low-level skill library that includes both MP algorithms and RL policies for efficient, adaptive, and high-precision behaviors in contact-rich assembly scenarios. Additionally, we design a high-level policy based on a Diffusion Transformer (DiT), trained via imitation learning with only a handful of demonstrations, to efficiently sequence and deploy these primitive skills.
  • We evaluate our framework on three long-horizon robotic assembly tasks, each involving the assembly of 4 to 9 objects. Experiments in both simulation and the real world show that, despite being trained on a single object, our approach generalizes to unseen objects and outperforms baseline methods in terms of both success rate and data efficiency.
2 Related Work
2.1 Task and Motion Planning
@@
Task and Motion Planning (TAMP) [19, 20, 21] is an effective approach for long-horizon manipulation problems as it can resolve temporally dependent constraints through hybrid symbolic-continuous reasoning [22, 23]. These plans may involve regrasping [24], clearing obstructing objects [25], or moving to gather information [26]. Despite these abilities, such methods typically require hand-designed symbolic transition functions and continuous parameter samplers. Additionally, TAMP methods are computationally expensive, which limits their ability to handle failures in dynamic tasks. We argue that the symbolic transition functions and samplers can be replaced by a learnable high-level policy to address the issues mentioned above.
2.2 Learning for Long-Horizon Manipulation
@@
There is a large body of research in learning for long-horizon manipulation tasks. Diffusion policies [9, 27, 28] have been shown to be a powerful tool for addressing complex manipulation tasks by leveraging their generative and multi-modal capabilities. However, they require a large amount of training data, i.e., human demonstrations, to learn effective end-to-end policies for complex, long-horizon tasks. Therefore, a significant amount of research has adopted the divide-and-conquer mindset and tackled the task by decomposing it into easier and reusable subtasks, instead of learning an entire task with a single policy.
@@
Skill decomposition and chaining [29, 30, 31, 32, 33, 34, 35, 36] is a promising way to synthesize long-horizon and complex behaviors by sequentially chaining previously learned simpler skills through transition functions. Mishra et al. [37] trains individual skill diffusion models as action primitives and combine them at test time, when learned distributions of the skills are linearly chained to solve for a long-horizon goal during evaluation. Mao et al. [38] learns primitive skills from human demonstrations based on haptic feedback and segments a long-horizon task into a sequence of skills for dexterous manipulation. Chen et al. [39] introduces a bi-directional optimization framework that chains RL-trained sub-policies together using a transition feasibility function for long-horizon dexterous manipulation.
2.3 Hierarchical Modeling in Robotics
@@
Hierarchical approaches typically offer policies at varying abstraction levels [40, 41, 42, 43, 44, 45]. Xian et al. [46] and Ma et al. [47] propose to learn a high-level policy from demonstrations to predict end-effector key poses and a low-level diffusion-based trajectory generator for connecting these key poses to achieve the final goal. MimicPlay [5] learns a high-level plan from human videos of manipulating objects and low-level visuomotor controls from teleoperated demonstrations on the real robot. MAPLE [48] enhances standard RL algorithms by incorporating a predefined library of behavioral primitives. The most relevant work to our approach uses hierarchical policy decomposition for a multi-stage cable routing task [49]. They define both scripted and imitation-learned low-level primitives, and train a high-level policy to select among them. In contrast, we incorporate both model-based and model-free low-level primitives, and learn a high-level policy that selects and instantiates parameterized primitives. Our RL policy for the contact-rich insertion primitive is trained in simulation and transferred to the real world via zero-shot sim-to-real deployment.
3 Method
Figure 2: We propose a hierarchical framework for long-horizon robotic assembly. The high-level policy $\pi_{\theta}$, trained via imitation learning, takes as input the object pose from a pose estimator and the robot’s proprioceptive state, and outputs a categorical distribution over low-level primitives and their parameters. The selected primitive is then executed by a low-level policy, which is either reinforcement learning (RL)-based or motion planning (MP)-based. Specifically, we train RL policies in simulation for contact-rich skills, such as ‘insert’, while MP policies are used for primitives operating in free space, such as ‘move’.
@@ 3 Method
In this work, we model the long-horizon robot assembly task as a parameterized-action Markov decision process (PAMDP) [50]. A PAMDP problem consists of the tuple $\langle\mathcal{S},\mathcal{A},P,R,\gamma\rangle$ where $\mathcal{S}$ is the continuous state space, $\mathcal{A}$ is a set of discrete action primitives $\{a_{1},a_{2},...,a_{k}\}$, each with $m_{a}$ parameters $X_{a}\subseteq{\mathbb{R}^{m_{a}}}$, $P(s,a,s^{\prime})$ is the probability of transitioning to state $s^{\prime}$ when taking action primitive $a$ in state $s$, $R(s,a)$ is the reward from taking action primitive $a$ in state $s$, and $\gamma$ is the discount factor. Each parameterized action primitive $a$ can be instantiated as either a model-based policy or an RL policy. Our goal in a PAMDP is to find a policy $\pi(a,x|s)$ that minimizes the discrepancy between the learned policy and the expert demonstrations that maximize reward under the PAMDP.
@@
Solving a PAMDP involves not only selecting the best discrete primitive but also optimizing over the parameter space for each primitive at each state. In this work, we fix a set of primitives that are building blocks of manipulation for assembly tasks (See Section 3.1) and attempt to learn the high-level policy $\pi_{\theta}(a,x|s)$ as a neural network with parameters $\theta$. This results in a hierarchical framework, shown in Figure 2. In this section, we describe our assumptions, the low-level primitive library, the high-level policy architecture, the data collection pipeline, and the pose estimation module.
We make the following assumptions in this work: (1) the CAD models of the assembly parts are known, as is common in industrial settings; (2) the base component of each assembly, as well as the fixture used for object reorientation, is fixed to the table. Under these assumptions, a single pose estimation step—combined with the known CAD models—is sufficient to determine the target insertion poses for each object with minimal error. Note that assumption (2) can be relaxed by performing continuous pose estimation, i.e., repeatedly estimating object poses throughout the task to account for potential shifts or disturbances.
@@
3.1 Low-level Primitive Library: Modeling Basic Skills with Motion Planning and Reinforcement Learning
@@
We focus on providing agents with a library of flexible primitives that act as foundational components for high-precision, contact-rich robotic assembly tasks. The hierarchical decision-making system operates independently of the specific implementations of these primitives, which may include either closed-loop, learning-based skills or model-based motion planners. Regardless of how they function internally, it is essential that the primitives are adaptable to varying behaviors, which is why we introduce parameters to customize each primitive. These parameters typically have clear semantics, such as an object index or a 6-DoF end-effector pose for a grasping primitive. While these primitives offer flexibility, we acknowledge that they are not exhaustive and can be extended.
[…]
@@
INSERT(o,g)(o,g): The robot has object $o$ in grasp and inserts it into its goal pose $g\in\mathbf{SE}(3)$. An RL policy $\pi_{L}(\phi)$ is employed for two primary reasons: first, insertion goal poses are subject to inaccuracies from pose estimation errors; second, insertion tasks inherently involve complex, contact-rich interactions between the object and its receptacle. We train an RL policy in simulation [51] using PPO [52]. Its observation space includes end-effector (EE), force-torque (FT), and EE pose relative to the goal pose; its action is EE velocity $u_{t}$; and its reward $\tilde{r}(s^{p}_{t},u_{t},g)$ is the negative distance between the current pose $s^{p}_{t}$ and the goal pose $g$.
@@
The objective of the goal-conditioned INSERT primitive is to reach the goal pose $g$ that maximizes the expectation of the cumulative return, as Equation 1 shows:
@@
3.2 High-level Policy: Composing Primitives via Imitation
As shown in Figure 2, the high-level policy $\pi_{\theta}(a,x|s)$ takes pose information obtained from pose estimation (Section 3.4) for a given object $o$ and robot proprioception as inputs $s$, to select the appropriate primitive $a$ from the low-level primitive library and instantiates it with parameters $x$ for low-level control. Parameters $x$ contains the object index $o$ and pose information obtained from pose estimation. We collect human demonstration data (Section 3.3) and train the high-level policy via imitation learning.
We modify the Diffusion Transformer (DiT) [53] architecture as the backbone of our high-level policy, considering that its strong sequential data handling capability is suitable for handling a history of previous states and actions. The DiT outputs softmax scores for each primitive and the continuous action parameter for the primitive with the highest score. These two functions are supervised using cross-entropy loss and MSE loss, respectively. We employ DiT blocks with adaptive LayerNorm-Zero conditioning. We use the Robomimic observation encoder [54] to extract features from object pose ($s^{op}$) and proprioceptive pose ($s^{p}$). The goal of imitation learning is to find a parameterized policy $\pi_{\theta}$ that can maximize the likelihood function based on the currently collected demonstration data $\mathcal{D}=\{(\mathbf{s,a,x)}\}$: